OpenStudy (anonymous):

List the coordinates of any y-intercept and (b) the coordinates of any x-intercept. Do not graph.

OpenStudy (phi):

you need to solve an equation to answer this. What is the equation ?

OpenStudy (anonymous):

y = x - 2; (3, 1) (-2, -4)

OpenStudy (phi):

List the coordinates of any y-intercept by definition, a y-intercept is where the line crosses the y axis when x is 0 to find the y-intercept, replace x with 0 in your equation. (That means, erase the x, and put 0 in its place): y = 0 - 2 0 -2 is just -2, so you get y= -2 the y-intercept will be at the point x=0, y= -2, written as (0,-2)

OpenStudy (phi):

for (b) the coordinates of any x-intercept. that means find x when y is 0: replace y with 0 in y = x - 2 what do you get for x ?

OpenStudy (anonymous):

That part I get, but how do I put this on a graph?

OpenStudy (phi):

the question says list the coordinates and Do not graph. Do you still want to graph this line ? or the x and y intercepts ?